calculate the temperature on the platinum scale corresponding to 400 degree celsius on a thermometer

1
Expert's answer
2021-11-08T11:52:33-0500

The temperature (in ohms) on the platinum thermometer scale:


"R(T)=R_0(1+a-\alpha T-\beta T^2),"

where

R0 - resistance at 0°C,

"a" - constant

"\alpha" - constant

β - constant.


Since the constants are not given, the resistance at 400°C can be found as


"R(400)=R_0(1+a-400\alpha-160000\beta)."

Plug in the missing values to get the answer.